OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge y = 2^{3^{x^2}}\]

OpenStudy (lgbasallote):

im thinking ln?

OpenStudy (lgbasallote):

\[\ln y = 3^{x^2} \ln 2\]

OpenStudy (lgbasallote):

\[\huge \frac 1y y' = 2x^2 3^{x^2} \ln 2\] is that right?

OpenStudy (lgbasallote):

\[\huge y' = (2^{3^{x^2}} )(2x^2)(3^{x^2})(\ln 2)\]

OpenStudy (lgbasallote):

did i do that right... \[a = 3^{x^2}\] \[\ln a = x^2 \ln 3\] \[\frac 1a a' = 2x \ln 3\] \[a' = 3^{x^2} 2x\ln 3\] oops

OpenStudy (alexwee123):

dy/dx= 2^(3x^2) * ln2 * ln 3 *2x? o.0

OpenStudy (lgbasallote):

my answer is \[\huge y' = (2^{3^{x^2}})(3^{x^2} 2x\ln 3)(\ln 2)\] that right?

OpenStudy (alexwee123):

thats what i got @lgbasallote :o

OpenStudy (lgbasallote):

haha yay...wonder if we're correct

OpenStudy (valpey):

me too \[y^\prime=(2^{3^{x^2}})(\ln{2})(3^{x^2})(\ln{3})2x\]

OpenStudy (lgbasallote):

ahh then we're all right since it's unanimous!!

OpenStudy (anonymous):

You are going right @lgbasallote